OpenStudy (anonymous):

Choose the correct simplification of the expression (5x^4y^2z^2)(3x^4y^3z^5). 15x^16y^6z^10 15x^8y^5z^7 8x^8y^5z^7 8x^16y^6z^10

OpenStudy (anonymous):

8x^8y^5z^7 my origional guess

OpenStudy (nurali):

Multiply cofficient that is 5*3=15 and add the power of same variables \[x^{4+4}=x^8, y^{2+3}=y^5, and z ^{2+5}=z^7\]

OpenStudy (anonymous):

thanks

OpenStudy (nurali):

My Pleasure.